OpenStudy (anonymous):

Fifty students in the third grade class listed their hair and eye colors in the table below: Brown hair Blonde hair Total Blue eyes 14 8 22 Brown eyes 16 12 28 39 20 50 Are the events "blonde hair" and "blue eyes" independent? Yes, P(blonde hair) • P(blue eyes) = P(blonde hair ∩ blue eyes) Yes, P(blonde hair) • P(blue eyes) ≠ P(blonde hair ∩ blue eyes) No, P(blonde hair) • P(blue eyes) = P(blonde hair ∩ blue eyes) No, P(blonde hair) • P(blue eyes) ≠ P(blonde hair ∩ blue eyes)

OpenStudy (anonymous):

@aum

OpenStudy (amistre64):

compare rows if they form a proportion, then we have independence

OpenStudy (amistre64):

comparing cols is the same, but i always do rows

OpenStudy (amistre64):

and its either yes, becuase they are equal or no becuase they isnt

OpenStudy (anonymous):

So what is the final answer

OpenStudy (amistre64):

please do not ask for a final answer, we approach a solution .. we dont just hand out solutions

OpenStudy (anonymous):

no I know the answer I just checking my work

OpenStudy (amistre64):

since we are comparing blond and blue: Brown hair Blonde hair Total Blue eyes 14 [8] 22 Brown eyes 16 12 28 39 20 50 does 8/22 = 20/50?

OpenStudy (amistre64):

50(8) = 22(20) 400 = not 400 looks like its no becuase not equal to me

OpenStudy (anonymous):

So the last answer choice correct as I chosen previously

OpenStudy (amistre64):

if im reading the tiny little print correctly, yes no and not equal

OpenStudy (anonymous):

So it is the last choice thanks